Fermi National Laboratory


Accelerator Update - June 9-June 11

Accelerator Overview

Monday June 9
The day shift began with D-zero (D0) experimenters accessing their collision hall, with the Antiproton Source (Pbar) stacking, and with MiniBooNE taking beam.

D0 came out of access at 8:31 AM.

At 9:23 AM, Operations put the Tevatron (TeV) into shot setup. Operations established store 2667 at 11:55 AM with an initial luminosity of 27.45E30.

The TeV front end crashed. A controls expert rebooted the computer at 3:12 PM.

At 10:38 PM, a Main Injector (MI) expert began conducting studies.

Tuesday June 10
The midnight shift began with TeV experts conducting end-of-store studies, with Pbar stacking, and with beam to MiniBooNE.

At 1:41 AM, Operations terminated store 2667. D0 accessed their collision hall.

D0 came out of access at 5:30 AM.

At 6:15 AM, TeV experts began a period of beam maintenance focusing on the sextupole magnets.

Operations put the TeV through a dry squeeze at 1:24 PM and then put the TeV into shot setup at 1:51 PM.

D0 reported having some trouble with their toroid magnet at 2:12 PM. An expert had the toroid working normally forty minutes later.

A Linac expert reported that Linac technicians were changing the driver tube for a RF station (LRF2). This held off all downstream beam.

At 3:30 PM, Operations established store 2671 with an initial luminosity of 28.04E30.

Linac techs completed their work on LRF2 at 4:57 AM. Pbar resumed stacking and MiniBooNE resumed taking beam.

The Collider Detector Facility (CDF) called complaining about the Tune of the beam. Operators made a horizontal change, which resolved the problem.

Wednesday June 11
The midnight shift began with Operations monitoring store 2671, with Pbar stacking, and with MiniBooNE taking beam.

At 3 AM, Operations worked on the Booster to resolve beam problems for MiniBooNE. Operators completed their tuning at 4 PM. MiniBooNE beam was off during this time.

Operations terminated store 2671 at 6:13 AM. TeV experts began a beam maintenance period.

A Switchyard expert began a study period at 6:23 AM. The expert reported getting beam to TeV sector F49 on the first pulse. This sector is the point where beam leaves the TeV tunnel and enters the Switchyard transfer lines.

Today's Plan
TeV experts are conducting aperture scans and other beam maintenance checks throughout the day. Operations will put the TeV into shot setup later this afternoon.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 6/13/2003   email Fermilab